Arrive at the Lewis structures of sncl5- and IF4+
1
Expert's answer
2020-02-17T07:06:48-0500

Solution.

"SnCl5^-"

"e = 4 + (5*7) + 1 = 40"





"IF4^+"

"e = 7 + (4*7) - 1 = 34"




Answer:

"SnCl5^-"


"IF4^+"